OpenStudy (anonymous):

if f(x)=x^2-x and g (x) = x-1, what is f(g(x))

OpenStudy (mathmate):

f(x)=x^2-x and g (x) = x-1 Start with f(x)=x^2-x replace all appearance of x with g(x)=(x-1): f(g(x)) = (x-1)^2-(x-1) Expand and get your answer.